OpenStudy (anonymous):

If A (0,1), B (2,4) and C (4,7) then B is the midpoint of AC? ture or false

OpenStudy (anonymous):

midpoint of A and C is given by x = (Ax+Cx)/2 , y = (Ay+Cy)/2 can you answer it now ?